OpenStudy (anonymous):

a line passes through A(3,7) and B(-4,9). Find the value of a if C (a,1) is on the line.

OpenStudy (anonymous):

Equation of the line is (9-7)/(-4-3) = (7-1)/(3-a) 2/-7=6/(3-a) 3-a=-7*6/2 a=24